Find the slope intercept equation of the horizontal line passing through the point (-5,-7): Find the slope intercept equation of the horizontal line passing through the point (-5,-7)
Answer by Fombitz(1152)

Slope-intercept form
y=mx+b
Slope of horizontal line is 0.
y=mx+b
y=0*x+b
y=b
Use your point (-5,-7):
y=b
-7=b
b=-7
Your equation is therefore,
y=-7